Suppose that you have borrowed $275,000 in the form of a 25-year loan with an annual interest rate of 5.5% with monthly payments and monthly compounding. How much principal will you pay in the 13th year of the loan?
Answer #1

Borrowed sum of money = 275,000

Total payments = 25*12 = 300

Interest rate per month = 5.5%/12 = 0.45833%

Monthly payment = 275000 * (A/P, 0.45833%, 300)

= 275000*(0.0045833*1.0045833^300)/(1.0045833^300 – 1)

= 1688.74

Total principal payment in 13th year = balance after 144th payment – balance after 156th payment

1688.74(P/A, 0.45833%, 300 - 144) – 1688.74(P/A, 0.45833%, 300 - 156)

= 187913.70 – 177729.90

= 10183.80

This is the required answer.